# Position IDs Matter: An Enhanced Position Layout for Efficient Context Compression in Large Language Models

### v3

@@ Abstract
Using special tokens (e.g., gist, memory, or compressed tokens) to compress context information is a common practice for large language models (LLMs). However, existing approaches often neglect that position encodings inherently induce local inductive biases in models, causing the compression process to ignore holistic contextual dependencies. We propose Enhanced Position Layout (EPL), a simple yet effective method that improves the context compression capability of LLMs by only adjusting position IDs, the numerical identifiers that specify token positions. EPL minimizes the distance between context tokens and their corresponding special tokens and at the same time maintains the sequence order in position IDs between context tokens, special tokens, and the subsequent tokens. Integrating EPL into our best performing context compression model results in 1.9 ROUGE-1 F1 improvement on out-of-domain question answering datasets in average. When extended to multimodal scenarios, EPL brings an average accuracy gain of 2.6 to vision compression LLMs.  [^1] We will release the code upon acceptance.
@@
Runsong Zhao1,*, Xin Liu1,*, Xinyu Liu1, Pengcheng Huang1, Chunyang Xiao, Tong Xiao1,2 and Jingbo Zhu1,2 1 NLP Lab, School of Computer Science and Engineering, Northeastern University, Shenyang, China 2 NiuTrans Research, Shenyang, China {fhe43694, chunyangx}@gmail.com {xiaotong, zhujingbo}@mail.neu.edu.com
11footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
In Transformer Vaswani et al. (2017) architectures, special tokens have been widely adopted as compression carriers of contextual information across natural language processing Devlin et al. (2019); Liu et al. (2019); Bulatov et al. (2022); Ge et al. (2024); Li et al. (2024b) and computer vision Dosovitskiy et al. (2021); Ye et al. (2025). For context compression, so-called soft prompt methods Chang et al. (2024); Li et al. (2025) employ encoders to condense long contexts into few special tokens, enabling decoders to perform inference based on compressed representations rather than raw inputs, thereby significantly reducing memory consumption and inference latency in long-context scenarios Jiang et al. (2024); Xu et al. (2024). We illustrate typical soft prompt architectures in Figure 1 where special tokens are appended at the end of the context, intending to capture the context semantics via causal attention mechanism in LLMs.
@@
The design ensures full context visibility for special tokens. However, we remark that in Transformer architectures position IDs do not need to coincide with physical token positions and the model’s perceived positional information is primarily determined by position IDs rather than physical token positions Vaswani et al. (2017). From this viewpoint, the local inductive biases introduced by position encodings Devlin et al. (2019); Vaswani et al. (2017); Su et al. (2023); Raffel et al. (2020); Press et al. (2022) weaken the efficacy of context compression under the default position layout (DPL), primarily because of the substantial distance between the special tokens and the context tokens, as illustrated in Figure 1 DPL. In this paper, we examine carefully position layout designs and propose Enhanced Position Layout (EPL) for soft prompt architectures, which comprises Uniform Position Layout (UPL) and Consistent Position Layout (CPL).
@@
UPL redistributes special tokens’ position IDs to achieve uniform distribution in the context tokens’ position ID space, as exemplified in Figure 1. By uniformly assigning position IDs amongst context token position IDs, a priori, most context tokens would have corresponding special tokens close to them. We assume that such a prior helps the special tokens compress the context. We formalize such intuitions, demonstrating the optimality of the UPL in Section 3.2.1. During compression, because special tokens are inserted and text chunks are reorganized, the position IDs between context, special tokens and subsequent tokens (e.g. reconstructed tokens or subsequent tokens such as QA pairs) can become inconsistent compared to their original positions before compression. Our proposed CPL in Section 3.2.2 guarantees to maintain the position ID sequence order for different tokens, between different text chunks in their natural causal order.
@@
We empirically apply EPL to two dominant context compression frameworks ICAE Ge et al. (2024) and 500xCompressor Li et al. (2024b). For the best model, on the autoencoding (AE) task, EPL yields a 1.8 BLEU gain and converges 9.7 times faster than DPL; on out-of-domain question answering (QA) tasks, EPL gives an average 1.9 ROUGE-1 F1 gain. When extending our application to multimodality with VoCo-LLaMA Ye et al. (2025), EPL yields an average 2.6 accuracy gain on multimodal benchmarks. The EPL improvement is consistent across base models of different scales. Further analysis shows that both UPL (which aims for better context compression) and CPL (which maintains causal sequence ordering) are essential for the final performance improvement across tasks. Finally, our UPL attention map visualization confirms the usefulness of our specified prior: UPL special tokens indeed focus more on tokens close to its assigned position IDs.
@@ 2.1 Local Bias of Position Encodings
Transformer architectures Vaswani et al. (2017) compute contextual token embeddings through position-invariant self-attention. Since natural language semantics crucially depend on token order, various position encodings (PEs) have been proposed to inject positional awareness including Sinusoidal PE Vaswani et al. (2017), RoPE Su et al. (2023), Learnable PE Devlin et al. (2019), T5 Bias Raffel et al. (2020) and ALiBi Press et al. (2022), etc. All approaches share the inductive bias that adjacent tokens should correlate more strongly. Taking the PEs with the trigonometric encoding (e.g. Sinusoidal/RoPE) design as examples, the position embedding at a certain position is mostly similar to its neighbors and the similarity decay as the distance increase, see Appendix H.1 for more details. ALiBi enforce such inductive bias by applying distance-sensitive penalties to attention scores. We further show in Section 5.3 that Learnable PE and T5-bias learn similar local bias through pre-training on natural text.
Figure 1: Comparison of UPL and DPL. In prior work (DPL), memory tokens are assigned position IDs 7 and 8. Our method (UPL) allocates them to position IDs 2 and 5. Tokens with ID in close proximity tend to exhibit higher attention scores.
@@ 2.2 Position Layout
This nuance is important but hardly noticeable because the default position layout (DPL) often coincides with the physical token positions as illustrated in DPL in Figure 1 (see Table 11 for more details on DPL). Such position layout is helpful for language modeling (LM) task since the task has been known to depend largely on its recent context Hu et al. (2024); Liu et al. (2024a). However, as LLMs are becoming a ubiquitous tool, we hypothesize that careful position layouts for some tasks can inject helpful inductive bias. In the rest of the paper, we focus on LLM compression tasks as our testbed.
@@ 3.1 Soft Prompt Methods
ICAE Ge et al. (2024) is a widely used encoder-decoder soft prompt method. Its encoder compresses long context into a few memory tokens, after which the decoder performs inference conditioned only on the memory tokens to achieve faster inference speed. Left of Figure 3 in Appendix illustrates this process through an example. ICAE can be trained through two stages: continued pretraining and fine-tuning. Continued pretraining trains on a combination of AutoEncoding (AE) tasks and Language Modeling (LM) tasks, which trains the LLM encoder so that the encoded memory tokens enable a frozen LLM to reconstruct losslessly the original context and at the same time predict the subsequent tokens following the context to maintain ICAE’s generation capability. Right of Figure 3 illustrates the AE training process. Fine-tuning further trains the ICAE encoder to adapt to real-word applications such as question answering (QA) that we use in this work. We review the AE and LM pretraining as well as QA finetuning in the next subsections. To handle arbitrarily long contexts, we adopt the multi-chunk version of ICAE, which divides any long context into chunks for independent compression and then aggregates the resulting memory tokens to represent the complete long context. 500xCompressor Ge et al. (2024) is similar to ICAE, with the difference of using the KV Cache of memory tokens as the compression carrier instead of the output of memory tokens.
@@ 3.1.1 Pretraining
During the pretraining stage, for a token sequence $X=\{x_{0},x_{1},\dots,x_{|X|-1}\}$, we take the first $p$ tokens as the context $X_{\text{context}}=\{x_{0},x_{1},\dots,x_{p-1}\}$ and the subsequent $|X|-p$ tokens as the completion $X_{\text{completion}}=\{x_{p},x_{p+1},\ldots,x_{|X|-1}\}$. The AE task only uses $X_{\text{context}}$, while the LM task leverages both $X_{\text{context}}$ and $X_{\text{completion}}$.
@@
where [;][;][ ; ] denotes concatenation along the sequence dimension. All $M^{(i)}$ share the same learnable parameters $M$, and $\theta_{\text{LoRA}}$ denotes a set of low-rank adapter Hu et al. (2022) parameters for the LLM. The final compressed representation is obtained by concatenating the results of each chunk: $\tilde{M}=[\tilde{M}^{(1)};\tilde{M}^{(2)};\dots;\tilde{M}^{(k)}]$ or $KV=[KV^{(1)};KV^{(2)};\dots;KV^{(k)}]$.
@@
[…]
where [AE] is a learnable token prompting the frozen decoder to generate $X_{\text{context}}$ as reconstruction. Similar to the AE task, the loss for LM task in ICAE is given by:
[…]
@@
EPL $1$ $2$ $510$ $3$ $8$ $508$ $511$ $512$ $1020$ $513$ $518$ $1018$
Table 1: Position layout of encoder for $S^{(1)}$ and $S^{(2)}$. More details see Appendix A, Table 12.
@@
EPL $3$ $508$ $513$ $1018$ 00 $1$ $2$ $1020$
Table 2: Position layout example of decoder in AE Task. More details see Appendix A, Table 13.
@@
EPL $3$ $508$ $513$ $1018$ $1020$ $1021$ $1022$ $2040$
Table 3: Position layout example of decoder in LM Task. More details see Appendix A, Table 14.
@@
[…]
@@ 3.1.2 Fine-tuning
ICAE Ge et al. (2024) allows further fine-tuning to enhance downstream task performance by enabling memory tokens to learn to focus on context that most relate to the task; the training process is similar to the LM pretraining. Let each training instance consists of a triplet $(C,Q,A)$, where context $C$ is compressed into either $\tilde{M}$ (ICAE) or $KV$ (500xCompressor). The answer $A=\{a_{0},a_{1},...,a_{|A|-1}\}$ is generated conditioned on the compressed representation and the question $Q=\{q_{0},q_{1},...,q_{|Q|-1}\}$. The loss for QA task in ICAE is given by:
@@ 3.2 Enhanced Position Layout
Memory tokens in soft prompt framework mainly aim to compress the context tokens so that the inference can be solely based on them to accelerate inference. However, the memory token DPL consists of a continuous range (i.e. $510$-$611$ in Table 1) and are all very distant from the context token’s range (i.e. 00-$509$). Given the local inductive bias of PEs that we briefly review in section 2.1, it would be advantageous to have memory token position IDs to be both close to context token position IDs and covering the context token ID range. In Section 3.2.1, we propose Uniform Position Layout (UPL) that has memory token position layout covering the context token ID range while achieving minimum ID distances between memory tokens and context tokens.
@@
[…]
@@ 3.2.1 Uniform Position Layout
3:$r\leftarrow(v_{L}-v_{1}+1)/|M|$
4:$o\leftarrow\frac{r-1}{2}$
5:$U^{*}\leftarrow\text{torch.linspace}\bigl{(}v_{1}+o,v_{L}-o,\;|M|\bigr{)}$
@@ 4.1 Experimental Setup
For continued pretraining, we utilize the SlimPajama-6B Soboleva et al. (2023) corpus. To evaluate model fine-tuning performance, we use MRQA Fisch et al. (2019a) dataset as our testbed. The dataset contains evaluation on both in-domain scenarios where the validation dataset has its training counterpart used during training and out-of-domain scenarios. We report results from both settings but mainly discuss results for out-of-domain scenarios as it assesses more critically the soft prompt compression effectiveness.
@@ 4.2 Fine-tuning Results
Following ICAE and 500xCompressor, we pretrain then fine-tune Llama-3.2-1B; we integrate our EPL changes described in Section 3.2 into different architectures respectively, then follow the same pretraining and finetuning steps. We evaluate the downstream performance using MRQA Fisch et al. (2019b) for different experiments. We assess the quality of the model’s answers using ROUGE-1 F1 Lin (2004) and Exact Match (EM) and report out-of-domain results in Table 4.
@@
Table 4: Pretraining and fine-tuning results. For the pretraining results, we report perplexity for completion and BLEU-4 Papineni et al. (2002) score for reconstruction calculated on a held-out set of 1k examples. For the fine-tuning results, we report out-of-domain results, which includes 6 datasets: BioASQ Tsatsaronis et al. (2015), DROP Dua et al. (2019), DuoRC Saha et al. (2018), RACE Lai et al. (2017), Relation Extraction (RE) Levy et al. (2017), and TextbookQA (TQA) Kembhavi et al. (2017). F1 in all the tables is ROUGE-1 F1 Lin (2004).
For both ICAE and 500xCompressor, incorporating EPL significantly improves the performance. The average ROUGE-1 F1 improves from 39.95 to 43.87 for ICAE and improves from 45.76 to 48.03 for 500xCompressor. The improvement was observed for most domains, suggesting that the method is overall effective. We observe a similar improvement for in-domain settings (see Table 16 in the appendix).

### v4

@@ Abstract
Using special tokens (e.g., gist, memory, or compressed tokens) to compress context information is a common practice for large language models (LLMs). However, existing approaches often neglect that position encodings inherently induce local inductive biases in models, causing the compression process to ignore holistic contextual dependencies. We propose Enhanced Position Layout (EPL), a simple yet effective method that improves the context compression capability of LLMs by only adjusting position IDs, the numerical identifiers that specify token positions. EPL minimizes the distance between context tokens and their corresponding special tokens and at the same time maintains the sequence order in position IDs between context tokens, special tokens, and the subsequent tokens. Integrating EPL into our best performing context compression model results in a 1.9 ROUGE-1 F1 improvement on out-of-domain question answering datasets on average. When extended to multimodal scenarios, EPL leads to an average accuracy gain of 2.6 points for vision compression LLMs.  [^1] Our code is available at https://github.com/1azybug/EPL.
@@
[…]
11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding author.
@@ 1 Introduction
In Transformer Vaswani et al. (2017) architectures, special tokens have been widely adopted as compression carriers of contextual information across natural language processing Devlin et al. (2019); Liu et al. (2019); Bulatov et al. (2022); Ge et al. (2024); Li et al. (2024b) and computer vision Dosovitskiy et al. (2021); Ye et al. (2025). For context compression, the so-called soft prompt methods Chang et al. (2024); Li et al. (2025) employ encoders to condense long contexts into few special tokens, enabling decoders to perform inference based on compressed representations rather than raw inputs, thereby significantly reducing memory consumption and inference latency in long-context scenarios Jiang et al. (2024); Xu et al. (2024); Luo et al. (2025). We illustrate typical soft prompt architectures in Figure 1 where special tokens are appended at the end of the context, in order to capture the context semantics via causal attention mechanism in LLMs.
@@
The design ensures full context visibility for special tokens. However, in Transformer architectures position IDs do not need to coincide with physical token positions and the model’s perceived positional information is primarily determined by position IDs rather than physical token positions Vaswani et al. (2017). From this viewpoint, the local inductive biases introduced by position encodings Devlin et al. (2019); Vaswani et al. (2017); Su et al. (2023); Raffel et al. (2020); Press et al. (2022) weaken the efficacy of context compression under the default position layout (DPL), due to the substantial distance between the special tokens and the context tokens, as shown in Figure 1 (DPL). In this paper, we carefully examine position layout designs and propose Enhanced Position Layout (EPL) for soft prompt architectures, which comprises Uniform Position Layout (UPL) and Consistent Position Layout (CPL).
@@
UPL redistributes special tokens’ position IDs to achieve uniform distribution in the context tokens’ position ID space, as exemplified in Figure 1. By uniformly assigning position IDs among context token position IDs, a priori, most context tokens would have corresponding special tokens close to them. We assume that such a prior helps the special tokens compress the context. We formalize such intuitions, demonstrating the optimality of the UPL in Section 3.2.1. During compression, because special tokens are inserted and text chunks are reorganized, the position IDs between context, special tokens and subsequent tokens (e.g. reconstructed tokens or subsequent tokens such as QA pairs) can become inconsistent compared to their original positions before compression. Our proposed CPL in Section 3.2.2 guarantees to maintain the order of the position ID sequence for different tokens, between different text chunks in their natural causal order.
@@
We empirically apply EPL to two dominant context compression frameworks: ICAE Ge et al. (2024) and 500xCompressor Li et al. (2024b). For the best model, on the autoencoding (AE) task, EPL yields a 1.8 BLEU gain and converges 9.7 times faster than DPL; on out-of-domain question answering (QA) tasks, EPL gives an average 1.9 ROUGE-1 F1 gain. To demonstrate the universality of EPL, we further conduct experiments by incorporating EPL into a more recent fine-grained compression framework Zhang et al. (2025) and test our EPL equipped compressor on PwC dataset Ge et al. (2024), an instruction-tuning dataset containing diverse context-related prompts. In both scenarios, incorporating EPL brings consistent improvement. When extending our application to multimodality with VoCo-LLaMA Ye et al. (2025), EPL yields an average 2.6 accuracy gain on multimodal benchmarks. The improvement of EPL is consistent across base models of different scales. Further analysis shows that both UPL (which aims for better context compression) and CPL (which maintains causal sequence ordering) are essential for the final performance improvement across tasks. Finally, our UPL attention map visualization confirms the usefulness of our specified prior: UPL special tokens indeed focus more on tokens close to its assigned position IDs.
@@ 2.1 Local Bias of Position Encodings
Transformer architectures Vaswani et al. (2017) compute contextual token embeddings through position-invariant self-attention. Since natural language semantics crucially depend on token order, various position encodings (PEs) have been proposed to inject positional awareness including Sinusoidal PE Vaswani et al. (2017), RoPE Su et al. (2023), Learnable PE Devlin et al. (2019), T5 Bias Raffel et al. (2020) and ALiBi Press et al. (2022), etc. All approaches share the inductive bias that adjacent tokens should correlate more strongly. Taking the PEs with the trigonometric encoding (e.g. Sinusoidal/RoPE) design as examples, the position embedding at a certain position is mostly similar to its neighbors and the similarity decays as the distance increases, see Appendix H.1 for more details. ALiBi enforces this inductive bias by applying distance-sensitive penalties to attention scores. We further show in Section 5.3 that Learnable PE and T5 Bias learn similar local bias through pre-training on natural text.
Figure 1: Comparison of UPL and DPL. In prior work (DPL), memory tokens are assigned position IDs 7 and 8. Our method (UPL) allocates them to position IDs 2 and 5. Tokens with IDs in close proximity tend to exhibit higher attention scores.
@@ 2.2 Position Layout
This nuance is important but hardly noticeable because the default position layout (DPL) often coincides with the physical token positions as illustrated in DPL in Figure 1 (see Table 14 for more details on DPL). Such position layout is helpful for language modeling (LM) task since the task depends largely on its recent context Hu et al. (2024); Liu et al. (2024a). However, as LLMs are becoming a ubiquitous tool, we hypothesize that careful position layouts for some tasks can inject helpful inductive bias. In the rest of the paper, we focus on LLM compression tasks as our testbed.
@@ 3.1 Soft Prompt Methods
ICAE Ge et al. (2024) is a widely used encoder-decoder soft prompt method. Its encoder compresses long context into a few memory tokens, after which the decoder performs inference conditioned only on the memory tokens to achieve faster inference speed. The left side of Figure 3 in the Appendix illustrates this process through an example. ICAE can be trained through two stages: continued pretraining and fine-tuning. Continued pretraining trains on a combination of AutoEncoding (AE) tasks and Language Modeling (LM) tasks, which trains the LLM encoder so that the encoded memory tokens enable a frozen LLM to losslessly reconstruct the original context and at the same time predict the subsequent tokens following the context to maintain ICAE’s generation capability. The right side of Figure 3 illustrates the AE training process. Fine-tuning further trains the ICAE encoder to adapt to real-world applications such as question answering (QA) that we use in this work. We review the AE and LM pretraining as well as QA fine-tuning in the next subsections. To handle arbitrarily long contexts, we adopt the multi-chunk version of ICAE, which divides any long context into chunks for independent compression and then aggregates the resulting memory tokens to represent the complete long context. 500xCompressor Ge et al. (2024) is similar to ICAE, with the difference of using the KV Cache of memory tokens as the compression carrier instead of the output of memory tokens.
@@ 3.1.1 Pretraining
During the pretraining stage, for a token sequence $X=\{x_{0},x_{1},\dots,x_{|X|-1}\}$, we take the first $p$ tokens as the context $X_{\text{context}}=\{x_{0},x_{1},\dots,x_{p-1}\}$ and the subsequent |X|p|X|-p tokens as the completion $X_{\text{completion}}=\{x_{p},x_{p+1},\ldots,x_{|X|-1}\}$. The AE task only uses $X_{\text{context}}$, while the LM task leverages both $X_{\text{context}}$ and $X_{\text{completion}}$.
@@
where $[;]$ denotes concatenation along the sequence dimension. All $M^{(i)}$ share the same learnable parameters $M$, and $\theta_{\text{LoRA}}$ denotes a set of low-rank adapter Hu et al. (2022) parameters for the LLM. The final compressed representation is obtained by concatenating the results of each chunk: $\tilde{M}=[\tilde{M}^{(1)};\tilde{M}^{(2)};\dots;\tilde{M}^{(k)}]$ or $KV=[KV^{(1)};KV^{(2)};\dots;KV^{(k)}]$.
@@
$\mathcal{L}_{\text{AE}}=-\log P\big(X_{\text{context}}\mid[\tilde{M};\texttt{[AE]}]\big)$ (2)
where [AE] is a learnable token prompting the frozen decoder to generate $X_{\text{context}}$ as reconstruction. Similar to the AE task, the loss for the LM task in ICAE is given by:
$\mathcal{L}_{\text{LM}}=-\log P\big(X_{\text{completion}}\mid[\tilde{M};\texttt{[LM]}]\big)$ (3)
@@
EPL 11 22 $510$ 33 88 $508$ $511$ $512$ $1020$ $513$ $518$ $1018$
Table 1: Position layout of the encoder for $S^{(1)}$ and $S^{(2)}$. We show DPL’s problematic position IDs in red, EPL’s UPL-adjusted IDs in green, and CPL-adjusted IDs in blue. For more details, see Appendix A, Table 15.
@@
EPL 33 $508$ $513$ $1018$ 0 11 22 $1020$
Table 2: Position layout example of the decoder in AE Task. For more details, see Appendix A, Table 16.
@@
EPL 33 $508$ $513$ $1018$ $1020$ $1021$ $1022$ $2040$
Table 3: Position layout example of the decoder in LM Task. For more details, see Appendix A, Table 17.
@@
The ICAE default position layout (DPL) always coincides with their physical token positions. For ICAE encoder for example, this means its DPL starts with position ID 0 and ranges till $|S^{(i)}|+|M^{(i)}|-1$. For 500xCompressor, its encoder DPL also coincides with physical token positions. However, the position IDs of the decoder’s $KV^{(i)}$ are the same as the position IDs of the encoder’s $M^{(i)}$ (the memory tokens of the i-th chunk) [^3] This is because the KV Cache has already cached the position information of the key_state, see this code snippet., which implies that the decoder DPL for the KVKV [^4] Recall that $KV$ concatenates all $KV^{(i)}$ with $KV=[KV^{(1)};KV^{(2)};\dots;KV^{(k)}]$. consists of $k$ repeated ranges from $|S^{(i)}|$ to $|S^{(i)}|+|M^{(i)}|-1$. During decoding, DPL for the rest of the tokens (e.g. $\texttt{[AE]},X_{\text{context}}$ in AE task) still coincides with their physical token positions. Table 2 and Table 3 show examples of ICAE and 500xCompressor DPL with $X_{\text{context}}$ having two chunks (kk=2) under LM and AE task respectively.
@@ 3.1.2 Fine-tuning
ICAE Ge et al. (2024) allows further fine-tuning to enhance downstream task performance by enabling memory tokens to learn to focus on the tokens that are most related to the task; the training process is similar to LM pretraining. Let each training instance consist of a triplet $(C,Q,A)$, where context $C$ is compressed into either $\tilde{M}$ (ICAE) or $KV$ (500xCompressor). The answer $A=\{a_{0},a_{1},...,a_{|A|-1}\}$ is generated conditioned on the compressed representation and the question $Q=\{q_{0},q_{1},...,q_{|Q|-1}\}$. The loss for QA task in ICAE is given by:
@@ 3.2 Enhanced Position Layout
Memory tokens in the soft prompt framework mainly aim to compress the context tokens so that the inference can be solely based on them to accelerate inference. However, the memory tokens in DPL consist of a continuous range (i.e. 510510-$611$ in Table 1) and are all very distant from the context token’s range (i.e. 0-$509$). Given the local inductive bias of PEs that we briefly review in Section 2.1, it would be advantageous to have memory token position IDs to be both close to context token position IDs and covering the context token ID range. In Section 3.2.1, we propose Uniform Position Layout (UPL) that has memory token position layout covering the context token ID range while achieving minimum ID distances between memory tokens and context tokens.
@@
Standard Transformer DPL coincides with physical token positions, which implies that tokens with larger position IDs follow the tokens with smaller position IDs, reflecting the causal relationship between the tokens through their assigned IDs. However, we observe that some DPL does not comply with such properties. For example, 500xCompressor’s decoder DPL as shown in Table 3 starts with the $KV$ position IDs $\{510$, 511, …, 611}611\} but is followed by position ID sequence $\{204,205,\dots,1224\}$ representing $\texttt{[LM]};X_{\text{completion}}]$. In Section 3.2.2, we detail our Consistent Position Layout (CPL) design which guarantees the resulting position layout to maintain the causal structure amongst position IDs. We hypothesize that aligning the causal structures for position IDs will benefit performance.
@@ 4.1 Experimental Setup
For continued pretraining, we utilize the SlimPajama-6B Soboleva et al. (2023) corpus. To evaluate model fine-tuning performance, we use the MRQA Fisch et al. (2019a) dataset as our testbed. The dataset contains evaluation on both in-domain scenarios where the validation dataset has its training counterpart used during training and out-of-domain scenarios. We report results from both settings but mainly discuss results for out-of-domain scenarios as it assesses more critically the soft prompt compression effectiveness.
@@ 4.2 Fine-tuning Results
Following ICAE and 500xCompressor, we pretrain then fine-tune Llama-3.2-1B; we integrate our EPL changes described in Section 3.2 into different architectures respectively, then follow the same pretraining and fine-tuning steps. We evaluate the downstream performance using MRQA Fisch et al. (2019b) for different experiments. We assess the quality of the model’s answers using ROUGE-1 F1 Lin (2004) and Exact Match (EM) and report out-of-domain results in Table 4.
@@
Table 4: Pretraining and fine-tuning results. For the pretraining results, we report perplexity for completion and BLEU-4 Papineni et al. (2002) score for reconstruction calculated on a held-out set of 1k examples. For the fine-tuning results, we report out-of-domain results, which include 6 datasets: BioASQ Tsatsaronis et al. (2015), DROP Dua et al. (2019), DuoRC Saha et al. (2018), RACE Lai et al. (2017), Relation Extraction (RE) Levy et al. (2017), and TextbookQA (TQA) Kembhavi et al. (2017). F1 in all the tables is ROUGE-1 F1 Lin (2004).
For both ICAE and 500xCompressor, incorporating EPL significantly improves the performance. The average ROUGE-1 F1 improves from 39.95 to 43.87 for ICAE and improves from 45.76 to 48.03 for 500xCompressor. The improvement is observed for most domains, suggesting that the method is overall effective. We observe a similar improvement for in-domain settings (see Table 19 in the appendix).
To further validate the effectiveness of EPL for general context compression, we test on the PwC dataset Ge et al. (2023) which contains diverse scenarios and question types. The results in Table 5 show that EPL significantly improves downstream task performance compared to its counterpart, demonstrating its effectiveness.
ROUGE-1 F1 BLEU4
ICAE-1B (DPL) 42.71 18.05
ICAE-1B (EPL) 45.59 21.04
Table 5: Results on PwC Ge et al. (2024).